### v1

# The Super-Kamiokande Collaboration

††thanks: also at BMCC/CUNY, Science Department, New York, New York, U.S.A.
The Super-Kamiokande Collaboration
††thanks: also at Université Paris-Saclay††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at Moscow Institute of Physics and Technology (MIPT), Moscow region, Russia and National Research Nuclear University ”MEPhI”, Moscow, Russia††thanks: also at IPSA-DRII, France††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at JINR, Dubna, Russia††thanks: also at the Graduate University of Science and Technology, Vietnam Academy of Science and Technology††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at Nambu Yoichiro Institute of Theoretical and Experimental Physics (NITEP)††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at Nambu Yoichiro Institute of Theoretical and Experimental Physics (NITEP)††thanks: also at BMCC/CUNY, Science Department, New York, New York, U.S.A.††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan
The T2K Collaboration
First joint oscillation analysis of Super-Kamiokande atmospheric and T2K accelerator neutrino data
K. Abe
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
S. Abe
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
C. Bronner
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
Y. Hayato
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
K. Hiraide
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
K. Hosokawa
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
K. Ieki
M. Ikeda
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
J. Kameda
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
Y. Kanemura
R. Kaneshima
Y. Kashiwagi
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
Y. Kataoka
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
S. Miki
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
S. Mine
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
University of California, Irvine, Department of Physics and Astronomy, Irvine, California, U.S.A.
M. Miura
S. Moriyama
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
M. Nakahata
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
Y. Nakano
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
S. Nakayama
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
Y. Noguchi
K. Sato
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
H. Sekiya
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
H. Shiba
K. Shimizu
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
M. Shiozawa
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
Y. Sonoda
Y. Suzuki
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
A. Takeda
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
Y. Takemoto
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
H. Tanaka
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
T. Yano
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
S. Han
University of Tokyo, Institute for Cosmic Ray Research, Research Center for Cosmic Neutrinos, Kashiwa, Japan
T. Kajita
University of Tokyo, Institute for Cosmic Ray Research, Research Center for Cosmic Neutrinos, Kashiwa, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
ILANCE, CNRS – University of Tokyo International Research Laboratory, Kashiwa, Chiba, Japan
K. Okumura
University of Tokyo, Institute for Cosmic Ray Research, Research Center for Cosmic Neutrinos, Kashiwa, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
T. Tashiro
T. Tomiya
X. Wang
S. Yoshida
University of Tokyo, Institute for Cosmic Ray Research, Research Center for Cosmic Neutrinos, Kashiwa, Japan
P. Fernandez
L. Labarga
N. Ospina
B. Zaldivar
University Autonoma Madrid, Department of Theoretical Physics, Madrid, Spain
B. W. Pointon
British Columbia Institute of Technology, Department of Physics, Burnaby, British Columbia, Canada
TRIUMF, Vancouver, British Columbia, Canada
E. Kearns
Boston University, Department of Physics, Boston, Massachusetts, U.S.A.
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
J. Mirabito
Boston University, Department of Physics, Boston, Massachusetts, U.S.A.
J. L. Raaf
Boston University, Department of Physics, Boston, Massachusetts, U.S.A.
L. Wan
Boston University, Department of Physics, Boston, Massachusetts, U.S.A.
T. Wester
Boston University, Department of Physics, Boston, Massachusetts, U.S.A.
J. Bian
N. J. Griskevich
University of California, Irvine, Department of Physics and Astronomy, Irvine, California, U.S.A.
M. B. Smy
H. W. Sobel
University of California, Irvine, Department of Physics and Astronomy, Irvine, California, U.S.A.
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
V. Takhistov
University of California, Irvine, Department of Physics and Astronomy, Irvine, California, U.S.A.
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
A. Yankelevich
University of California, Irvine, Department of Physics and Astronomy, Irvine, California, U.S.A.
J. Hill
California State University, Department of Physics, Dominguez Hills, Carson, California, U.S.A
M. C. Jang
S. H. Lee
D. H. Moon
R. G. Park
Chonnam National University, Institute for Universe and Elementary Particles, Gwangju, Korea
B. Bodur
Duke University, Department of Physics, Durham, North Carolina, U.S.A.
K. Scholberg
C. W. Walter
Duke University, Department of Physics, Durham, North Carolina, U.S.A.
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
A. Beauchêne
O. Drapier
A. Giampaolo
Th. A. Mueller
A. D. Santos
P. Paganini
B. Quilain
R. Rogly
Ecole Polytechnique, IN2P3-CNRS, Laboratoire Leprince-Ringuet, Palaiseau, France
T. Nakamura
Gifu University, Department of Physics, Gifu, Japan
J. S. Jang
Gwangju Institute of Science and Technology, GIST College, Gwangju, Korea
L. N. Machado
University of Glasgow, School of Physics and Astronomy, Glasgow, United Kingdom
J. G. Learned
University of Hawaii, Department of Physics and Astronomy, Honolulu, Hawaii, U.S.A.
K. Choi
N. Iovine
Institute for Basic Science (IBS), Center for Underground Physics, Daejeon, Korea
S. Cao
Institute For Interdisciplinary Research in Science and Education (IFIRSE), ICISE, Quy Nhon, Vietnam
L. H. V. Anthony
D. Martin
N. W. Prouse
M. Scott
Y. Uchida
Imperial College London, Department of Physics, London, United Kingdom
V. Berardi
N. F. Calabria
M. G. Catanesi
E. Radicioni
INFN Sezione di Bari and Università e Politecnico di Bari, Dipartimento Interuniversitario di Fisica, Bari, Italy
A. Langella
G. De Rosa
INFN Sezione di Napoli and Università di Napoli, Dipartimento di Fisica, Napoli, Italy
G. Collazuol
M. Feltre
F. Iacob
M. Mattiazzi
INFN Sezione di Padova and Università di Padova, Dipartimento di Fisica, Padova, Italy
L. Ludovici
INFN Sezione di Roma and Università di Roma “La Sapienza”, Roma, Italy
M. Gonin
L. Périssé
G. Pronost
ILANCE, CNRS – University of Tokyo International Research Laboratory, Kashiwa, Chiba, Japan
C. Fujisawa
S. Horiuchi
M. Kobayashi
Y.M. Liu
Y. Maekawa
Y. Nishimura
R. Okazaki
Keio University, Department of Physics, Kanagawa, Japan
R. Akutsu
M. Friend
T. Hasegawa
T. Ishida
T. Kobayashi
M. Jakkapu
T. Matsubara
T. Nakadaira
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
K. Nakamura
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
Y. Oyama
A. Portocarrero Yrey
K. Sakashita
T. Sekiguchi
T. Tsukamoto
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
N. Bhuiyan
G. T. Burton
F. Di Lodovico
J. Gao
A. Goldsack
T. Katori
J. Migenda
R. M. Ramsden
Z. Xie
King’s College London, Department of Physics, Strand, London, United Kingdom
S. Zsoldos
King’s College London, Department of Physics, Strand, London, United Kingdom
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
A. T. Suzuki
Y. Takagi
Kobe University, Kobe, Japan
Y. Takeuchi
Kobe University, Kobe, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
H. Zhong
Kobe University, Kobe, Japan
J. Feng
L. Feng
J. R. Hu
Z. Hu
M. Kawaue
T. Kikawa
M. Mori
Kyoto University, Department of Physics, Kyoto, Japan
T. Nakaya
Kyoto University, Department of Physics, Kyoto, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
R. A. Wendell
Kyoto University, Department of Physics, Kyoto, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
K. Yasutome
Kyoto University, Department of Physics, Kyoto, Japan
S. J. Jenkins
N. McCauley
P. Mehta
A. Tarrant
University of Liverpool, Department of Physics, Liverpool, United Kingdom
M. J. Wilking
University of Minnesota, School of Physics and Astronomy, Minneapolis, Minnesota, U.S.A.
Y. Fukuda
Miyagi University of Education, Department of Physics, Sendai, Japan
Y. Itow
Nagoya University, Institute for Space-Earth Environmental Research, Nagoya, Aichi, Japan
Nagoya University, Kobayashi-Maskawa Institute for the Origin of Particles and the Universe, Nagoya, Aichi, Japan
H. Menjo
K. Ninomiya
Y. Yoshioka
Nagoya University, Institute for Space-Earth Environmental Research, Nagoya, Aichi, Japan
J. Lagoda
M. Mandal
P. Mijakowski
Y. S. Prabhu
J. Zalipska
National Centre for Nuclear Research, Warsaw, Poland
M. Jia
J. Jiang
W. Shi
C. Yanagisawa
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
M. Harada
Y. Hino
H. Ishino
Okayama University, Department of Physics, Okayama, Japan
Y. Koshio
Okayama University, Department of Physics, Okayama, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
F. Nakanishi
S. Sakai
T. Tada
T. Tano
Okayama University, Department of Physics, Okayama, Japan
T. Ishizuka
Osaka Electro-Communication University, Media Communication Center, Neyagawa, Osaka, Japan
G. Barr
D. Barrow
Oxford University, Department of Physics, Oxford, United Kingdom
L. Cook
Oxford University, Department of Physics, Oxford, United Kingdom
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
S. Samani
Oxford University, Department of Physics, Oxford, United Kingdom
D. Wark
Oxford University, Department of Physics, Oxford, United Kingdom
STFC, Rutherford Appleton Laboratory, Harwell Oxford, and Daresbury Laboratory, Warrington, United Kingdom
A. Holin
F. Nova
Rutherford Appleton Laboratory, Harwell, Oxford, United Kingdom
S. Jung
B. S. Yang
J. Y. Yang
J. Yoo
Seoul National University, Department of Physics, Seoul, Korea
J. E. P. Fannon
L. Kneale
M. Malek
J. M. McElwee
M. D. Thiesse
L. F. Thompson
S. T. Wilson
University of Sheffield, Department of Physics and Astronomy, Sheffield, United Kingdom
H. Okazawa
Shizuoka University of Welfare, Department of Informatics in Social Welfare, Yaizu, Shizuoka, Japan
S. M. Lakshmi
University of Silesia, Institute of Physics, Katowice, Poland
S. B. Kim
E. Kwon
J. W. Seo
I. Yu
Sungkyunkwan University, Department of Physics, Suwon, Korea
A. K. Ichikawa
K. D. Nakamura
S. Tairafune
Tohoku University, Faculty of Science, Department of Physics, Miyagi, Japan
K. Nishijima
Tokai University, Department of Physics, Hiratsuka, Kanagawa, Japan
A. Eguchi
K. Nakagiri
University of Tokyo, Department of Physics, Tokyo, Japan
Y. Nakajima
University of Tokyo, Department of Physics, Tokyo, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
S. Shima
N. Taniuchi
E. Watanabe
University of Tokyo, Department of Physics, Tokyo, Japan
M. Yokoyama
University of Tokyo, Department of Physics, Tokyo, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
P. de Perio
S. Fujita
C. Jesús-Valls
K. Martens
K. M. Tsui
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
M. R. Vagins
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
University of California, Irvine, Department of Physics and Astronomy, Irvine, California, U.S.A.
J. Xia
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
S. Izumiyama
M. Kuze
R. Matsumoto
K. Terada
Tokyo Institute of Technology, Department of Physics, Tokyo, Japan
R. Asaka
M. Ishitsuka
H. Ito
Y. Ommura
N. Shigeta
M. Shinoki
K. Yamauchi
T. Yoshida
Tokyo University of Science, Faculty of Science and Technology, Department of Physics, Noda, Chiba, Japan
R. Gaur
TRIUMF, Vancouver, British Columbia, Canada
V. Gousy-Leblanc
also at University of Victoria, Department of Physics and Astronomy, PO Box 1700 STN CSC, Victoria, BC V8W 2Y2, Canada.
TRIUMF, Vancouver, British Columbia, Canada
M. Hartz
A. Konaka
X. Li
TRIUMF, Vancouver, British Columbia, Canada
S. Chen
B. D. Xu
Y. Wu
A.Q. Zhang
B. Zhang
Tsinghua University, Department of Engineering Physics, Beijing, China
M. Posiadala-Zezula
University of Warsaw, Faculty of Physics, Warsaw, Poland
S. B. Boyd
R. Edwards
D. Hadley
M. Nicholson
M. O’Flaherty
B. Richards
University of Warwick, Department of Physics, Coventry, United Kingdom
A. Ali
University of Winnipeg, Department of Physics, Winnipeg, Manitoba, Canada
TRIUMF, Vancouver, British Columbia, Canada
B. Jamieson
University of Winnipeg, Department of Physics, Winnipeg, Manitoba, Canada
S. Amanai
Ll. Marti
A. Minamino
R. Shibayama
R. Shimamura
S. Suzuki
Yokohama National University, Department of Physics, Yokohama, Japan
University Autonoma Madrid, Department of Theoretical Physics, Madrid, Spain
University of Bern, Albert Einstein Center for Fundamental Physics, Laboratory for High Energy Physics (LHEP), Bern, Switzerland
Boston University, Department of Physics, Boston, Massachusetts, U.S.A.
University of British Columbia, Department of Physics and Astronomy, Vancouver, British Columbia, Canada
University of California, Irvine, Department of Physics and Astronomy, Irvine, California, U.S.A.
IRFU, CEA, Université Paris-Saclay, Gif-sur-Yvette, France
University of Colorado at Boulder, Department of Physics, Boulder, Colorado, U.S.A.
Colorado State University, Department of Physics, Fort Collins, Colorado, U.S.A.
Duke University, Department of Physics, Durham, North Carolina, U.S.A.
Eötvös Loránd University, Department of Atomic Physics, Budapest, Hungary
ETH Zurich, Institute for Particle Physics and Astrophysics, Zurich, Switzerland
Hanoi University of Science, Vietnam National University, Hanoi, Vietnam
CERN European Organization for Nuclear Research, Genéve, Switzerland
University of Geneva, Section de Physique, DPNC, Geneva, Switzerland
University of Glasgow, School of Physics and Astronomy, Glasgow, United Kingdom
Ghent University, Department of Physics and Astronomy, Gent, Belgium
H. Niewodniczanski Institute of Nuclear Physics PAN, Cracow, Poland
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
University of Houston, Department of Physics, Houston, Texas, U.S.A.
Institut de Fisica d’Altes Energies (IFAE) - The Barcelona Institute of Science and Technology, Campus UAB, Bellaterra (Barcelona) Spain
Institut für Physik, Johannes Gutenberg-Universität Mainz, Staudingerweg 7, Mainz, Germany
IFIC (CSIC & University of Valencia), Valencia, Spain
Institute For Interdisciplinary Research in Science and Education (IFIRSE), ICISE, Quy Nhon, Vietnam
Imperial College London, Department of Physics, London, United Kingdom
INFN Sezione di Bari and Università e Politecnico di Bari, Dipartimento Interuniversitario di Fisica, Bari, Italy
INFN Sezione di Napoli and Università di Napoli, Dipartimento di Fisica, Napoli, Italy
INFN Sezione di Padova and Università di Padova, Dipartimento di Fisica, Padova, Italy
INFN Sezione di Roma and Università di Roma “La Sapienza”, Roma, Italy
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
International Centre of Physics, Institute of Physics (IOP), Vietnam Academy of Science and Technology (VAST), 10 Dao Tan, Ba Dinh, Hanoi, Vietnam
ILANCE, CNRS – University of Tokyo International Research Laboratory, Kashiwa, Chiba, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
Keio University, Department of Physics, Kanagawa, Japan
King’s College London, Department of Physics, Strand, London, United Kingdom
Kobe University, Kobe, Japan
Kyoto University, Department of Physics, Kyoto, Japan
Lancaster University, Physics Department, Lancaster, United Kingdom
Lawrence Berkeley National Laboratory, Berkeley, California, U.S.A.
Ecole Polytechnique, IN2P3-CNRS, Laboratoire Leprince-Ringuet, Palaiseau, France
University of Liverpool, Department of Physics, Liverpool, United Kingdom
Louisiana State University, Department of Physics and Astronomy, Baton Rouge, Louisiana, U.S.A.
Joint Institute for Nuclear Research, Dubna, Moscow Region, Russia
Michigan State University, Department of Physics and Astronomy, East Lansing, Michigan, U.S.A.
Miyagi University of Education, Department of Physics, Sendai, Japan
National Centre for Nuclear Research, Warsaw, Poland
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
STFC, Rutherford Appleton Laboratory, Harwell Oxford, and Daresbury Laboratory, Warrington, United Kingdom
Okayama University, Department of Physics, Okayama, Japan
Osaka Metropolitan University, Department of Physics, Osaka, Japan
Oxford University, Department of Physics, Oxford, United Kingdom
University of Pennsylvania, Department of Physics and Astronomy, Philadelphia, Pennsylvania U.S.A.
University of Pittsburgh, Department of Physics and Astronomy, Pittsburgh, Pennsylvania, U.S.A.
Queen Mary University of London, School of Physics and Astronomy, London, United Kingdom
University of Regina, Department of Physics, Regina, Saskatchewan, Canada
University of Rochester, Department of Physics and Astronomy, Rochester, New York, U.S.A.
Royal Holloway University of London, Department of Physics, Egham, Surrey, United Kingdom
RWTH Aachen University, III. Physikalisches Institut, Aachen, Germany
University of Minnesota, School of Physics and Astronomy, Minneapolis, Minnesota, U.S.A.
Universidad de Sevilla, Departamento de Física Atómica, Molecular y Nuclear, Sevilla, Spain
University of Sheffield, Department of Physics and Astronomy, Sheffield, United Kingdom
University of Silesia, Institute of Physics, Katowice, Poland
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
South Dakota School of Mines and Technology, Rapid City, South Dakota, U.S.A.
University of Tokyo, Department of Physics, Tokyo, Japan
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
University of Tokyo, Institute for Cosmic Ray Research, Research Center for Cosmic Neutrinos, Kashiwa, Japan
Tokyo Institute of Technology, Department of Physics, Tokyo, Japan
Tokyo Metropolitan University, Department of Physics, Tokyo, Japan
Tokyo University of Science, Faculty of Science and Technology, Department of Physics, Noda, Chiba, Japan
University of Toronto, Department of Physics, Toronto, Ontario, Canada
TRIUMF, Vancouver, British Columbia, Canada
University of Warsaw, Faculty of Physics, Warsaw, Poland
Warsaw University of Technology, Institute of Radioelectronics and Multimedia Technology, Warsaw, Poland
Tohoku University, Faculty of Science, Department of Physics, Miyagi, Japan
University of Warwick, Department of Physics, Coventry, United Kingdom
University of Winnipeg, Department of Physics, Winnipeg, Manitoba, Canada
Wroclaw University, Faculty of Physics and Astronomy, Wroclaw, Poland
Yokohama National University, Department of Physics, Yokohama, Japan
York University, Department of Physics and Astronomy, Toronto, Ontario, Canada
K. Abe
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
S. Abe
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
N. Akhlaq
Queen Mary University of London, School of Physics and Astronomy, London, United Kingdom
R. Akutsu
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
H. Alarakia-Charles
Lancaster University, Physics Department, Lancaster, United Kingdom
A. Ali
University of Winnipeg, Department of Physics, Winnipeg, Manitoba, Canada
TRIUMF, Vancouver, British Columbia, Canada
Y.I. Alj Hakim
Imperial College London, Department of Physics, London, United Kingdom
S. Alonso Monsalve
ETH Zurich, Institute for Particle Physics and Astrophysics, Zurich, Switzerland
C. Andreopoulos
University of Liverpool, Department of Physics, Liverpool, United Kingdom
L. Anthony
Imperial College London, Department of Physics, London, United Kingdom
M. Antonova
IFIC (CSIC & University of Valencia), Valencia, Spain
S. Aoki
Kobe University, Kobe, Japan
K.A. Apte
Imperial College London, Department of Physics, London, United Kingdom
T. Arai
University of Tokyo, Department of Physics, Tokyo, Japan
T. Arihara
Tokyo Metropolitan University, Department of Physics, Tokyo, Japan
S. Arimoto
Kyoto University, Department of Physics, Kyoto, Japan
Y. Asada
Yokohama National University, Department of Physics, Yokohama, Japan
Y. Ashida
Kyoto University, Department of Physics, Kyoto, Japan
E.T. Atkin
Imperial College London, Department of Physics, London, United Kingdom
N. Babu
Louisiana State University, Department of Physics and Astronomy, Baton Rouge, Louisiana, U.S.A.
M. Barbi
University of Regina, Department of Physics, Regina, Saskatchewan, Canada
G.J. Barker
University of Warwick, Department of Physics, Coventry, United Kingdom
G. Barr
Oxford University, Department of Physics, Oxford, United Kingdom
D. Barrow
Oxford University, Department of Physics, Oxford, United Kingdom
P. Bates
University of Liverpool, Department of Physics, Liverpool, United Kingdom
M. Batkiewicz-Kwasniak
H. Niewodniczanski Institute of Nuclear Physics PAN, Cracow, Poland
V. Berardi
INFN Sezione di Bari and Università e Politecnico di Bari, Dipartimento Interuniversitario di Fisica, Bari, Italy
L. Berns
Tohoku University, Faculty of Science, Department of Physics, Miyagi, Japan
S. Bhadra
York University, Department of Physics and Astronomy, Toronto, Ontario, Canada
A. Blanchet
University of Geneva, Section de Physique, DPNC, Geneva, Switzerland
A. Blondel
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
University of Geneva, Section de Physique, DPNC, Geneva, Switzerland
S. Bolognesi
IRFU, CEA, Université Paris-Saclay, Gif-sur-Yvette, France
S. Bordoni
University of Geneva, Section de Physique, DPNC, Geneva, Switzerland
S.B. Boyd
University of Warwick, Department of Physics, Coventry, United Kingdom
A. Bravar
University of Geneva, Section de Physique, DPNC, Geneva, Switzerland
C. Bronner
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
A. Bubak
University of Silesia, Institute of Physics, Katowice, Poland
M. Buizza Avanzini
Ecole Polytechnique, IN2P3-CNRS, Laboratoire Leprince-Ringuet, Palaiseau, France
J.A. Caballero
Universidad de Sevilla, Departamento de Física Atómica, Molecular y Nuclear, Sevilla, Spain
N.F. Calabria
INFN Sezione di Bari and Università e Politecnico di Bari, Dipartimento Interuniversitario di Fisica, Bari, Italy
S. Cao
Institute For Interdisciplinary Research in Science and Education (IFIRSE), ICISE, Quy Nhon, Vietnam
D. Carabadjac
Ecole Polytechnique, IN2P3-CNRS, Laboratoire Leprince-Ringuet, Palaiseau, France
A.J. Carter
Royal Holloway University of London, Department of Physics, Egham, Surrey, United Kingdom
S.L. Cartwright
University of Sheffield, Department of Physics and Astronomy, Sheffield, United Kingdom
M.P. Casado
Institut de Fisica d’Altes Energies (IFAE) - The Barcelona Institute of Science and Technology, Campus UAB, Bellaterra (Barcelona) Spain
M.G. Catanesi
INFN Sezione di Bari and Università e Politecnico di Bari, Dipartimento Interuniversitario di Fisica, Bari, Italy
A. Cervera
IFIC (CSIC & University of Valencia), Valencia, Spain
J. Chakrani
Lawrence Berkeley National Laboratory, Berkeley, California, U.S.A.
A. Chalumeau
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
D. Cherdack
University of Houston, Department of Physics, Houston, Texas, U.S.A.
P.S. Chong
University of Pennsylvania, Department of Physics and Astronomy, Philadelphia, Pennsylvania U.S.A.
A. Chvirova
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
M. Cicerchia
INFN Sezione di Padova and Università di Padova, Dipartimento di Fisica, Padova, Italy
J. Coleman
University of Liverpool, Department of Physics, Liverpool, United Kingdom
G. Collazuol
INFN Sezione di Padova and Università di Padova, Dipartimento di Fisica, Padova, Italy
L. Cook
Oxford University, Department of Physics, Oxford, United Kingdom
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
F. Cormier
TRIUMF, Vancouver, British Columbia, Canada
A. Cudd
University of Colorado at Boulder, Department of Physics, Boulder, Colorado, U.S.A.
D. D’ago
INFN Sezione di Padova and Università di Padova, Dipartimento di Fisica, Padova, Italy
C. Dalmazzone
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
T. Daret
IRFU, CEA, Université Paris-Saclay, Gif-sur-Yvette, France
P. Dasgupta
Eötvös Loránd University, Department of Atomic Physics, Budapest, Hungary
C. Davis
University of Pennsylvania, Department of Physics and Astronomy, Philadelphia, Pennsylvania U.S.A.
Yu.I. Davydov
Joint Institute for Nuclear Research, Dubna, Moscow Region, Russia
A. De Roeck
CERN European Organization for Nuclear Research, Genéve, Switzerland
G. De Rosa
INFN Sezione di Napoli and Università di Napoli, Dipartimento di Fisica, Napoli, Italy
T. Dealtry
Lancaster University, Physics Department, Lancaster, United Kingdom
C.C. Delogu
INFN Sezione di Padova and Università di Padova, Dipartimento di Fisica, Padova, Italy
C. Densham
STFC, Rutherford Appleton Laboratory, Harwell Oxford, and Daresbury Laboratory, Warrington, United Kingdom
A. Dergacheva
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
R. Dharmapal
Wroclaw University, Faculty of Physics and Astronomy, Wroclaw, Poland
F. Di Lodovico
King’s College London, Department of Physics, Strand, London, United Kingdom
G. Diaz Lopez
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
S. Dolan
CERN European Organization for Nuclear Research, Genéve, Switzerland
D. Douqa
University of Geneva, Section de Physique, DPNC, Geneva, Switzerland
T.A. Doyle
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
O. Drapier
Ecole Polytechnique, IN2P3-CNRS, Laboratoire Leprince-Ringuet, Palaiseau, France
K.E. Duffy
Oxford University, Department of Physics, Oxford, United Kingdom
J. Dumarchez
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
P. Dunne
Imperial College London, Department of Physics, London, United Kingdom
K. Dygnarowicz
Warsaw University of Technology, Institute of Radioelectronics and Multimedia Technology, Warsaw, Poland
A. Eguchi
University of Tokyo, Department of Physics, Tokyo, Japan
J. Elias
University of Rochester, Department of Physics and Astronomy, Rochester, New York, U.S.A.
S. Emery-Schrenk
IRFU, CEA, Université Paris-Saclay, Gif-sur-Yvette, France
G. Erofeev
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
A. Ershova
Ecole Polytechnique, IN2P3-CNRS, Laboratoire Leprince-Ringuet, Palaiseau, France
G. Eurin
IRFU, CEA, Université Paris-Saclay, Gif-sur-Yvette, France
D. Fedorova
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
S. Fedotov
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
M. Feltre
INFN Sezione di Padova and Università di Padova, Dipartimento di Fisica, Padova, Italy
L. Feng
Kyoto University, Department of Physics, Kyoto, Japan
D. Ferlewicz
University of Tokyo, Department of Physics, Tokyo, Japan
A.J. Finch
Lancaster University, Physics Department, Lancaster, United Kingdom
G.A. Fiorentini Aguirre
York University, Department of Physics and Astronomy, Toronto, Ontario, Canada
G. Fiorillo
INFN Sezione di Napoli and Università di Napoli, Dipartimento di Fisica, Napoli, Italy
M.D. Fitton
STFC, Rutherford Appleton Laboratory, Harwell Oxford, and Daresbury Laboratory, Warrington, United Kingdom
J.M. Franco Patiño
Universidad de Sevilla, Departamento de Física Atómica, Molecular y Nuclear, Sevilla, Spain
M. Friend
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
Y. Fujii
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
Y. Fukuda
Miyagi University of Education, Department of Physics, Sendai, Japan
Y. Furui
Tokyo Metropolitan University, Department of Physics, Tokyo, Japan
L. Giannessi
University of Geneva, Section de Physique, DPNC, Geneva, Switzerland
C. Giganti
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
V. Glagolev
Joint Institute for Nuclear Research, Dubna, Moscow Region, Russia
M. Gonin
ILANCE, CNRS – University of Tokyo International Research Laboratory, Kashiwa, Chiba, Japan
J. González Rosa
Universidad de Sevilla, Departamento de Física Atómica, Molecular y Nuclear, Sevilla, Spain
E.A.G. Goodman
University of Glasgow, School of Physics and Astronomy, Glasgow, United Kingdom
A. Gorin
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
K. Gorshanov
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
M. Grassi
INFN Sezione di Padova and Università di Padova, Dipartimento di Fisica, Padova, Italy
M. Guigue
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
D.R. Hadley
University of Warwick, Department of Physics, Coventry, United Kingdom
J.T. Haigh
University of Warwick, Department of Physics, Coventry, United Kingdom
S. Han
University of Tokyo, Institute for Cosmic Ray Research, Research Center for Cosmic Neutrinos, Kashiwa, Japan
D.A. Harris
York University, Department of Physics and Astronomy, Toronto, Ontario, Canada
M. Hartz
TRIUMF, Vancouver, British Columbia, Canada
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
T. Hasegawa
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
S. Hassani
IRFU, CEA, Université Paris-Saclay, Gif-sur-Yvette, France
N.C. Hastings
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
Y. Hayato
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
I. Heitkamp
Tohoku University, Faculty of Science, Department of Physics, Miyagi, Japan
D. Henaff
INFN Sezione di Padova and Università di Padova, Dipartimento di Fisica, Padova, Italy
Y. Hino
Okayama University, Department of Physics, Okayama, Japan
M. Hogan
Colorado State University, Department of Physics, Fort Collins, Colorado, U.S.A.
J. Holeczek
University of Silesia, Institute of Physics, Katowice, Poland
A. Holin
STFC, Rutherford Appleton Laboratory, Harwell Oxford, and Daresbury Laboratory, Warrington, United Kingdom
T. Holvey
Oxford University, Department of Physics, Oxford, United Kingdom
N.T. Hong Van
International Centre of Physics, Institute of Physics (IOP), Vietnam Academy of Science and Technology (VAST), 10 Dao Tan, Ba Dinh, Hanoi, Vietnam
T. Honjo
Osaka Metropolitan University, Department of Physics, Osaka, Japan
K. Hosokawa
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
J. Hu
Kyoto University, Department of Physics, Kyoto, Japan
A.K. Ichikawa
Tohoku University, Faculty of Science, Department of Physics, Miyagi, Japan
K. Ieki
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
M. Ikeda
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
T. Ishida
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
M. Ishitsuka
Tokyo University of Science, Faculty of Science and Technology, Department of Physics, Noda, Chiba, Japan
A. Izmaylov
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
M. Jakkapu
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
B. Jamieson
University of Winnipeg, Department of Physics, Winnipeg, Manitoba, Canada
S.J. Jenkins
University of Liverpool, Department of Physics, Liverpool, United Kingdom
C. Jesús-Valls
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
M. Jia
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
J.J. Jiang
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
J.Y. Ji
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
P. Jonsson
Imperial College London, Department of Physics, London, United Kingdom
S. Joshi
IRFU, CEA, Université Paris-Saclay, Gif-sur-Yvette, France
C.K. Jung
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
M. Kabirnezhad
Imperial College London, Department of Physics, London, United Kingdom
A.C. Kaboth
Royal Holloway University of London, Department of Physics, Egham, Surrey, United Kingdom
STFC, Rutherford Appleton Laboratory, Harwell Oxford, and Daresbury Laboratory, Warrington, United Kingdom
T. Kajita
University of Tokyo, Institute for Cosmic Ray Research, Research Center for Cosmic Neutrinos, Kashiwa, Japan
H. Kakuno
Tokyo Metropolitan University, Department of Physics, Tokyo, Japan
J. Kameda
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
S. Karpova
University of Geneva, Section de Physique, DPNC, Geneva, Switzerland
S.P. Kasetti
Louisiana State University, Department of Physics and Astronomy, Baton Rouge, Louisiana, U.S.A.
V.S. Kasturi
University of Geneva, Section de Physique, DPNC, Geneva, Switzerland
Y. Kataoka
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
T. Katori
King’s College London, Department of Physics, Strand, London, United Kingdom
Y. Kawamura
Osaka Metropolitan University, Department of Physics, Osaka, Japan
M. Kawaue
Kyoto University, Department of Physics, Kyoto, Japan
E. Kearns
Boston University, Department of Physics, Boston, Massachusetts, U.S.A.
M. Khabibullin
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
A. Khotjantsev
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
T. Kikawa
Kyoto University, Department of Physics, Kyoto, Japan
S. King
King’s College London, Department of Physics, Strand, London, United Kingdom
V. Kiseeva
Joint Institute for Nuclear Research, Dubna, Moscow Region, Russia
J. Kisiel
University of Silesia, Institute of Physics, Katowice, Poland
L. Kneale
University of Sheffield, Department of Physics and Astronomy, Sheffield, United Kingdom
H. Kobayashi
University of Tokyo, Department of Physics, Tokyo, Japan
T. Kobayashi
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
L. Koch
Institut für Physik, Johannes Gutenberg-Universität Mainz, Staudingerweg 7, Mainz, Germany
S. Kodama
University of Tokyo, Department of Physics, Tokyo, Japan
M. Kolupanova
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
A. Konaka
TRIUMF, Vancouver, British Columbia, Canada
L.L. Kormos
Lancaster University, Physics Department, Lancaster, United Kingdom
Y. Koshio
Okayama University, Department of Physics, Okayama, Japan
T. Koto
Tokyo Metropolitan University, Department of Physics, Tokyo, Japan
K. Kowalik
National Centre for Nuclear Research, Warsaw, Poland
Y. Kudenko
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
Y. Kudo
Yokohama National University, Department of Physics, Yokohama, Japan
S. Kuribayashi
Kyoto University, Department of Physics, Kyoto, Japan
R. Kurjata
Warsaw University of Technology, Institute of Radioelectronics and Multimedia Technology, Warsaw, Poland
V. Kurochka
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
T. Kutter
Louisiana State University, Department of Physics and Astronomy, Baton Rouge, Louisiana, U.S.A.
M. Kuze
Tokyo Institute of Technology, Department of Physics, Tokyo, Japan
M. La Commara
INFN Sezione di Napoli and Università di Napoli, Dipartimento di Fisica, Napoli, Italy
L. Labarga
University Autonoma Madrid, Department of Theoretical Physics, Madrid, Spain
M. Lachat
University of Rochester, Department of Physics and Astronomy, Rochester, New York, U.S.A.
K. Lachner
University of Warwick, Department of Physics, Coventry, United Kingdom
J. Lagoda
National Centre for Nuclear Research, Warsaw, Poland
S.M. Lakshmi
University of Silesia, Institute of Physics, Katowice, Poland
M. Lamers James
Lancaster University, Physics Department, Lancaster, United Kingdom
STFC, Rutherford Appleton Laboratory, Harwell Oxford, and Daresbury Laboratory, Warrington, United Kingdom
A. Langella
INFN Sezione di Napoli and Università di Napoli, Dipartimento di Fisica, Napoli, Italy
J.-F. Laporte
IRFU, CEA, Université Paris-Saclay, Gif-sur-Yvette, France
D. Last
University of Rochester, Department of Physics and Astronomy, Rochester, New York, U.S.A.
N. Latham
University of Warwick, Department of Physics, Coventry, United Kingdom
M. Laveder
INFN Sezione di Padova and Università di Padova, Dipartimento di Fisica, Padova, Italy
L. Lavitola
INFN Sezione di Napoli and Università di Napoli, Dipartimento di Fisica, Napoli, Italy
M. Lawe
Lancaster University, Physics Department, Lancaster, United Kingdom
Y. Lee
Kyoto University, Department of Physics, Kyoto, Japan
D. Leon Silverio
South Dakota School of Mines and Technology, Rapid City, South Dakota, U.S.A.
S. Levorato
INFN Sezione di Padova and Università di Padova, Dipartimento di Fisica, Padova, Italy
S. Lewis
King’s College London, Department of Physics, Strand, London, United Kingdom
C. Lin
Imperial College London, Department of Physics, London, United Kingdom
R.P. Litchfield
University of Glasgow, School of Physics and Astronomy, Glasgow, United Kingdom
S.L. Liu
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
W. Li
Oxford University, Department of Physics, Oxford, United Kingdom
A. Longhin
INFN Sezione di Padova and Università di Padova, Dipartimento di Fisica, Padova, Italy
K.R. Long
Imperial College London, Department of Physics, London, United Kingdom
STFC, Rutherford Appleton Laboratory, Harwell Oxford, and Daresbury Laboratory, Warrington, United Kingdom
A. Lopez Moreno
King’s College London, Department of Physics, Strand, London, United Kingdom
L. Ludovici
INFN Sezione di Roma and Università di Roma “La Sapienza”, Roma, Italy
X. Lu
University of Warwick, Department of Physics, Coventry, United Kingdom
T. Lux
Institut de Fisica d’Altes Energies (IFAE) - The Barcelona Institute of Science and Technology, Campus UAB, Bellaterra (Barcelona) Spain
L.N. Machado
University of Glasgow, School of Physics and Astronomy, Glasgow, United Kingdom
L. Magaletti
INFN Sezione di Bari and Università e Politecnico di Bari, Dipartimento Interuniversitario di Fisica, Bari, Italy
K. Mahn
Michigan State University, Department of Physics and Astronomy, East Lansing, Michigan, U.S.A.
K.K. Mahtani
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
M. Malek
University of Sheffield, Department of Physics and Astronomy, Sheffield, United Kingdom
M. Mandal
National Centre for Nuclear Research, Warsaw, Poland
S. Manly
University of Rochester, Department of Physics and Astronomy, Rochester, New York, U.S.A.
A.D. Marino
University of Colorado at Boulder, Department of Physics, Boulder, Colorado, U.S.A.
L. Marti-Magro
Yokohama National University, Department of Physics, Yokohama, Japan
D.G.R. Martin
Imperial College London, Department of Physics, London, United Kingdom
M. Martini
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
J.F. Martin
University of Toronto, Department of Physics, Toronto, Ontario, Canada
T. Maruyama
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
T. Matsubara
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
R. Matsumoto
Tokyo Institute of Technology, Department of Physics, Tokyo, Japan
V. Matveev
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
C. Mauger
University of Pennsylvania, Department of Physics and Astronomy, Philadelphia, Pennsylvania U.S.A.
K. Mavrokoridis
University of Liverpool, Department of Physics, Liverpool, United Kingdom
E. Mazzucato
IRFU, CEA, Université Paris-Saclay, Gif-sur-Yvette, France
N. McCauley
University of Liverpool, Department of Physics, Liverpool, United Kingdom
K.S. McFarland
University of Rochester, Department of Physics and Astronomy, Rochester, New York, U.S.A.
C. McGrew
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
J. McKean
Imperial College London, Department of Physics, London, United Kingdom
A. Mefodiev
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
G.D. Megias
Universidad de Sevilla, Departamento de Física Atómica, Molecular y Nuclear, Sevilla, Spain
P. Mehta
University of Liverpool, Department of Physics, Liverpool, United Kingdom
L. Mellet
Michigan State University, Department of Physics and Astronomy, East Lansing, Michigan, U.S.A.
C. Metelko
University of Liverpool, Department of Physics, Liverpool, United Kingdom
M. Mezzetto
INFN Sezione di Padova and Università di Padova, Dipartimento di Fisica, Padova, Italy
S. Miki
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
E. Miller
King’s College London, Department of Physics, Strand, London, United Kingdom
A. Minamino
Yokohama National University, Department of Physics, Yokohama, Japan
O. Mineev
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
S. Mine
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
University of California, Irvine, Department of Physics and Astronomy, Irvine, California, U.S.A.
M. Miura
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
L. Molina Bueno
IFIC (CSIC & University of Valencia), Valencia, Spain
S. Moriyama
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
S. Moriyama
Yokohama National University, Department of Physics, Yokohama, Japan
P. Morrison
University of Glasgow, School of Physics and Astronomy, Glasgow, United Kingdom
Th.A. Mueller
Ecole Polytechnique, IN2P3-CNRS, Laboratoire Leprince-Ringuet, Palaiseau, France
D. Munford
University of Houston, Department of Physics, Houston, Texas, U.S.A.
A. Muñoz
Ecole Polytechnique, IN2P3-CNRS, Laboratoire Leprince-Ringuet, Palaiseau, France
ILANCE, CNRS – University of Tokyo International Research Laboratory, Kashiwa, Chiba, Japan
L. Munteanu
CERN European Organization for Nuclear Research, Genéve, Switzerland
K. Nagai
Yokohama National University, Department of Physics, Yokohama, Japan
Y. Nagai
Eötvös Loránd University, Department of Atomic Physics, Budapest, Hungary
T. Nakadaira
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
K. Nakagiri
University of Tokyo, Department of Physics, Tokyo, Japan
M. Nakahata
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
Y. Nakajima
University of Tokyo, Department of Physics, Tokyo, Japan
A. Nakamura
Okayama University, Department of Physics, Okayama, Japan
K. Nakamura
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
K.D. Nakamura
Tohoku University, Faculty of Science, Department of Physics, Miyagi, Japan
Y. Nakano
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
S. Nakayama
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
T. Nakaya
Kyoto University, Department of Physics, Kyoto, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
K. Nakayoshi
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
C.E.R. Naseby
Imperial College London, Department of Physics, London, United Kingdom
T.V. Ngoc
Kyoto University, Department of Physics, Kyoto, Japan
D.T. Nguyen
Hanoi University of Science, Vietnam National University, Hanoi, Vietnam
V.Q. Nguyen
Ecole Polytechnique, IN2P3-CNRS, Laboratoire Leprince-Ringuet, Palaiseau, France
K. Niewczas
Ghent University, Department of Physics and Astronomy, Gent, Belgium
S. Nishimori
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
Y. Nishimura
Keio University, Department of Physics, Kanagawa, Japan
Y. Noguchi
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
T. Nosek
National Centre for Nuclear Research, Warsaw, Poland
F. Nova
STFC, Rutherford Appleton Laboratory, Harwell Oxford, and Daresbury Laboratory, Warrington, United Kingdom
P. Novella
IFIC (CSIC & University of Valencia), Valencia, Spain
J.C. Nugent
Tohoku University, Faculty of Science, Department of Physics, Miyagi, Japan
H.M. O’Keeffe
Lancaster University, Physics Department, Lancaster, United Kingdom
L. O’Sullivan
Institut für Physik, Johannes Gutenberg-Universität Mainz, Staudingerweg 7, Mainz, Germany
T. Odagawa
Kyoto University, Department of Physics, Kyoto, Japan
R. Okazaki
Keio University, Department of Physics, Kanagawa, Japan
W. Okinaga
University of Tokyo, Department of Physics, Tokyo, Japan
K. Okumura
University of Tokyo, Institute for Cosmic Ray Research, Research Center for Cosmic Neutrinos, Kashiwa, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
T. Okusawa
Osaka Metropolitan University, Department of Physics, Osaka, Japan
N. Onda
Kyoto University, Department of Physics, Kyoto, Japan
N. Ospina
University Autonoma Madrid, Department of Theoretical Physics, Madrid, Spain
L. Osu
Ecole Polytechnique, IN2P3-CNRS, Laboratoire Leprince-Ringuet, Palaiseau, France
Y. Oyama
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
V. Palladino
INFN Sezione di Napoli and Università di Napoli, Dipartimento di Fisica, Napoli, Italy
V. Paolone
University of Pittsburgh, Department of Physics and Astronomy, Pittsburgh, Pennsylvania, U.S.A.
M. Pari
INFN Sezione di Padova and Università di Padova, Dipartimento di Fisica, Padova, Italy
J. Parlone
University of Liverpool, Department of Physics, Liverpool, United Kingdom
J. Pasternak
Imperial College London, Department of Physics, London, United Kingdom
D. Payne
University of Liverpool, Department of Physics, Liverpool, United Kingdom
G.C. Penn
University of Liverpool, Department of Physics, Liverpool, United Kingdom
D. Pershey
Duke University, Department of Physics, Durham, North Carolina, U.S.A.
M. Pfaff
Imperial College London, Department of Physics, London, United Kingdom
L. Pickering
STFC, Rutherford Appleton Laboratory, Harwell Oxford, and Daresbury Laboratory, Warrington, United Kingdom
G. Pintaudi
Yokohama National University, Department of Physics, Yokohama, Japan
C. Pistillo
University of Bern, Albert Einstein Center for Fundamental Physics, Laboratory for High Energy Physics (LHEP), Bern, Switzerland
B. Popov
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
A.J. Portocarrero Yrey
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
K. Porwit
University of Silesia, Institute of Physics, Katowice, Poland
M. Posiadala-Zezula
University of Warsaw, Faculty of Physics, Warsaw, Poland
Y.S. Prabhu
National Centre for Nuclear Research, Warsaw, Poland
H. Prasad
Wroclaw University, Faculty of Physics and Astronomy, Wroclaw, Poland
F. Pupilli
INFN Sezione di Padova and Università di Padova, Dipartimento di Fisica, Padova, Italy
B. Quilain
Ecole Polytechnique, IN2P3-CNRS, Laboratoire Leprince-Ringuet, Palaiseau, France
P.T. Quyen
Institute For Interdisciplinary Research in Science and Education (IFIRSE), ICISE, Quy Nhon, Vietnam
T. Radermacher
RWTH Aachen University, III. Physikalisches Institut, Aachen, Germany
E. Radicioni
INFN Sezione di Bari and Università e Politecnico di Bari, Dipartimento Interuniversitario di Fisica, Bari, Italy
B. Radics
York University, Department of Physics and Astronomy, Toronto, Ontario, Canada
M.A. Ramirez
University of Pennsylvania, Department of Physics and Astronomy, Philadelphia, Pennsylvania U.S.A.
P.N. Ratoff
Lancaster University, Physics Department, Lancaster, United Kingdom
M. Reh
University of Colorado at Boulder, Department of Physics, Boulder, Colorado, U.S.A.
C. Riccio
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
E. Rondio
National Centre for Nuclear Research, Warsaw, Poland
S. Roth
RWTH Aachen University, III. Physikalisches Institut, Aachen, Germany
N. Roy
York University, Department of Physics and Astronomy, Toronto, Ontario, Canada
A. Rubbia
ETH Zurich, Institute for Particle Physics and Astrophysics, Zurich, Switzerland
L. Russo
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
A. Rychter
Warsaw University of Technology, Institute of Radioelectronics and Multimedia Technology, Warsaw, Poland
W. Saenz
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
K. Sakashita
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
F. Sánchez
University of Geneva, Section de Physique, DPNC, Geneva, Switzerland
Y. Sato
Tokyo University of Science, Faculty of Science and Technology, Department of Physics, Noda, Chiba, Japan
T. Schefke
Louisiana State University, Department of Physics and Astronomy, Baton Rouge, Louisiana, U.S.A.
C.M. Schloesser
University of Geneva, Section de Physique, DPNC, Geneva, Switzerland
K. Scholberg
Duke University, Department of Physics, Durham, North Carolina, U.S.A.
M. Scott
Imperial College London, Department of Physics, London, United Kingdom
Y. Seiya
Osaka Metropolitan University, Department of Physics, Osaka, Japan
T. Sekiguchi
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
H. Sekiya
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
D. Sgalaberna
ETH Zurich, Institute for Particle Physics and Astrophysics, Zurich, Switzerland
A. Shaikhiev
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
M. Shiozawa
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
Y. Shiraishi
Okayama University, Department of Physics, Okayama, Japan
A. Shvartsman
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
N. Skrobova
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
K. Skwarczynski
Royal Holloway University of London, Department of Physics, Egham, Surrey, United Kingdom
D. Smyczek
RWTH Aachen University, III. Physikalisches Institut, Aachen, Germany
M. Smy
University of California, Irvine, Department of Physics and Astronomy, Irvine, California, U.S.A.
J.T. Sobczyk
Wroclaw University, Faculty of Physics and Astronomy, Wroclaw, Poland
H. Sobel
University of California, Irvine, Department of Physics and Astronomy, Irvine, California, U.S.A.
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
F.J.P. Soler
University of Glasgow, School of Physics and Astronomy, Glasgow, United Kingdom
A.J. Speers
Lancaster University, Physics Department, Lancaster, United Kingdom
R. Spina
INFN Sezione di Bari and Università e Politecnico di Bari, Dipartimento Interuniversitario di Fisica, Bari, Italy
Y. Stroke
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
I.A. Suslov
Joint Institute for Nuclear Research, Dubna, Moscow Region, Russia
S. Suvorov
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
A. Suzuki
Kobe University, Kobe, Japan
S.Y. Suzuki
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
Y. Suzuki
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
M. Tada
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
S. Tairafune
Tohoku University, Faculty of Science, Department of Physics, Miyagi, Japan
A. Takeda
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
Y. Takeuchi
Kobe University, Kobe, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
K. Takifuji
Tohoku University, Faculty of Science, Department of Physics, Miyagi, Japan
H.K. Tanaka
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
H. Tanigawa
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
A. Teklu
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
V.V. Tereshchenko
Joint Institute for Nuclear Research, Dubna, Moscow Region, Russia
N. Thamm
RWTH Aachen University, III. Physikalisches Institut, Aachen, Germany
L.F. Thompson
University of Sheffield, Department of Physics and Astronomy, Sheffield, United Kingdom
W. Toki
Colorado State University, Department of Physics, Fort Collins, Colorado, U.S.A.
C. Touramanis
University of Liverpool, Department of Physics, Liverpool, United Kingdom
T. Tsukamoto
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
M. Tzanov
Louisiana State University, Department of Physics and Astronomy, Baton Rouge, Louisiana, U.S.A.
Y. Uchida
Imperial College London, Department of Physics, London, United Kingdom
M. Vagins
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
University of California, Irvine, Department of Physics and Astronomy, Irvine, California, U.S.A.
D. Vargas
Institut de Fisica d’Altes Energies (IFAE) - The Barcelona Institute of Science and Technology, Campus UAB, Bellaterra (Barcelona) Spain
M. Varghese
Institut de Fisica d’Altes Energies (IFAE) - The Barcelona Institute of Science and Technology, Campus UAB, Bellaterra (Barcelona) Spain
G. Vasseur
IRFU, CEA, Université Paris-Saclay, Gif-sur-Yvette, France
E. Villa
CERN European Organization for Nuclear Research, Genéve, Switzerland
University of Geneva, Section de Physique, DPNC, Geneva, Switzerland
W.G.S. Vinning
University of Warwick, Department of Physics, Coventry, United Kingdom
U. Virginet
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
T. Vladisavljevic
STFC, Rutherford Appleton Laboratory, Harwell Oxford, and Daresbury Laboratory, Warrington, United Kingdom
T. Wachala
H. Niewodniczanski Institute of Nuclear Physics PAN, Cracow, Poland
D. Wakabayashi
Tohoku University, Faculty of Science, Department of Physics, Miyagi, Japan
H.T. Wallace
University of Sheffield, Department of Physics and Astronomy, Sheffield, United Kingdom
J.G. Walsh
Michigan State University, Department of Physics and Astronomy, East Lansing, Michigan, U.S.A.
Y. Wang
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
L. Wan
Boston University, Department of Physics, Boston, Massachusetts, U.S.A.
D. Wark
STFC, Rutherford Appleton Laboratory, Harwell Oxford, and Daresbury Laboratory, Warrington, United Kingdom
Oxford University, Department of Physics, Oxford, United Kingdom
M.O. Wascko
Oxford University, Department of Physics, Oxford, United Kingdom
STFC, Rutherford Appleton Laboratory, Harwell Oxford, and Daresbury Laboratory, Warrington, United Kingdom
A. Weber
Institut für Physik, Johannes Gutenberg-Universität Mainz, Staudingerweg 7, Mainz, Germany
R. Wendell
Kyoto University, Department of Physics, Kyoto, Japan
M.J. Wilking
University of Minnesota, School of Physics and Astronomy, Minneapolis, Minnesota, U.S.A.
C. Wilkinson
Lawrence Berkeley National Laboratory, Berkeley, California, U.S.A.
J.R. Wilson
King’s College London, Department of Physics, Strand, London, United Kingdom
K. Wood
Lawrence Berkeley National Laboratory, Berkeley, California, U.S.A.
C. Wret
Oxford University, Department of Physics, Oxford, United Kingdom
J. Xia
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
Y.-h. Xu
Lancaster University, Physics Department, Lancaster, United Kingdom
K. Yamamoto
Osaka Metropolitan University, Department of Physics, Osaka, Japan
T. Yamamoto
Osaka Metropolitan University, Department of Physics, Osaka, Japan
C. Yanagisawa
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
G. Yang
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
T. Yano
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
K. Yasutome
Kyoto University, Department of Physics, Kyoto, Japan
N. Yershov
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
U. Yevarouskaya
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
M. Yokoyama
University of Tokyo, Department of Physics, Tokyo, Japan
Y. Yoshimoto
University of Tokyo, Department of Physics, Tokyo, Japan
N. Yoshimura
Kyoto University, Department of Physics, Kyoto, Japan
M. Yu
York University, Department of Physics and Astronomy, Toronto, Ontario, Canada
R. Zaki
York University, Department of Physics and Astronomy, Toronto, Ontario, Canada
A. Zalewska
H. Niewodniczanski Institute of Nuclear Physics PAN, Cracow, Poland
J. Zalipska
National Centre for Nuclear Research, Warsaw, Poland
K. Zaremba
Warsaw University of Technology, Institute of Radioelectronics and Multimedia Technology, Warsaw, Poland
G. Zarnecki
H. Niewodniczanski Institute of Nuclear Physics PAN, Cracow, Poland
J. Zhang
TRIUMF, Vancouver, British Columbia, Canada
University of British Columbia, Department of Physics and Astronomy, Vancouver, British Columbia, Canada
X.Y. Zhao
ETH Zurich, Institute for Particle Physics and Astrophysics, Zurich, Switzerland
H. Zhong
Kobe University, Kobe, Japan
T. Zhu
Imperial College London, Department of Physics, London, United Kingdom
M. Ziembicki
Warsaw University of Technology, Institute of Radioelectronics and Multimedia Technology, Warsaw, Poland
E.D. Zimmerman
University of Colorado at Boulder, Department of Physics, Boulder, Colorado, U.S.A.
M. Zito
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
S. Zsoldos
King’s College London, Department of Physics, Strand, London, United Kingdom
(May 24, 2024)
Abstract
The Super-Kamiokande and T2K collaborations present a joint measurement of neutrino oscillation parameters from their atmospheric and beam neutrino data. It uses a common interaction model for events overlapping in neutrino energy and correlated detector systematic uncertainties between the two datasets, which are found to be compatible. Using 3244.4 days of atmospheric data and a beam exposure of protons on target in (anti)neutrino mode, the analysis finds a 1.9 exclusion of CP-conservation (defined as ) and a preference for the normal mass ordering.
Introduction—Following the observation of neutrino oscillations [1], experiments now aim to fully characterize the three-flavor mixing paradigm described by the Pontecorvo–Maki–Nakagawa–Sakata matrix. Here, neutrino mixing is governed by three mixing angles (, , and ), two mass splittings ( and ), and one Charge Parity (CP) violating phase (). While some oscillation parameters have been precisely measured [2], others remain relatively unconstrained. In particular, the CP-violating phase, the ordering of the neutrino mass states (MO), and the octant of have not been determined experimentally. The Jarlskog invariant [3, 4],
, gives the magnitude of a potential CP violation.
Experimental setup—The Super-Kamiokande (SK) experiment [5] measures atmospheric neutrino oscillations using a large multi-purpose water Cherenkov detector located in the Kamioka mine in Gifu, Japan. The detector has a 32 kiloton inner detector optically separated from a 2 meter thick outer detector, which mainly serves as a veto region. Atmospheric neutrinos, produced by the interaction of cosmic rays with nuclei in the Earth’s atmosphere, include a mixture of neutrino flavor states, as well as a wide range of propagation baselines (1513000 km) and neutrino energies (MeVTeV).
The Tokai-to-Kamioka (T2K) long-baseline neutrino experiment [6] measures neutrino oscillations over a baseline of 295 km using a primarily muon (anti)neutrino beam produced by the neutrino facility at J-PARC, located in Ibaraki, Japan. SK is T2K’s far detector and measures neutrinos after oscillations 2.5∘ off of the beam axis. The beam neutrino flux and neutrino interaction cross sections are constrained by a suite of near detectors (T2K ND) situated 280 m downstream of the neutrino production target.
Motivation for a joint analysis—T2K’s off-axis neutrino beam provides a narrow energy spectrum peaked at 600 MeV and a known direction for beam-induced events at SK. This enables a precise determination of and using the “disappearance” channels (), and some constraint on the CP-violating phase and the MO via the lower statistics “appearance” channels (). The beam composition can be changed from primarily neutrinos to primarily antineutrinos by switching the polarity of electromagnetic focusing horns, providing the ability to compare neutrino and antineutrino oscillations.
T2K’s baseline probes the first oscillation maximum and matter effects have a limited impact. Nonetheless, there are notable parameter degeneracies between the MO and [7]. Degeneracies between the lower and upper octants of also affect the measurements.
SK’s atmospheric neutrino sample, on the other hand, provides a comparatively weak constraint on the atmospheric mixing parameters, due to limited information about the incoming neutrino direction, and a broader range of neutrino energies.
However, upward-going neutrinos experience large matter effects, producing asymmetric oscillations between neutrinos and antineutrinos and providing sensitivity to both the octant of and the MO through the appearance channel at 210 GeV.
This provides a means of breaking the MO– degeneracy, complementing the MO sensitivity achieved at T2K through its lower energy appearance events.
Additionally, common systematic uncertainties at SK can be better constrained in a joint fit than in the individual experiments, due to the addition of the T2K ND constraints to the SK analysis and the large statistics atmospheric sample to the T2K analysis.
Analysis strategy—The analysis described here is based on previous analyses from the two experiments [8, 9], modified to produce a coherent joint analysis. Neutrino oscillation parameters are measured by comparing predictions for the rates and spectra of the atmospheric and beam neutrinos to observations performed at SK and making statistical inferences, following both frequentist and Bayesian approaches. The predictions are made using a model of the two experiments, covering neutrino fluxes, interactions, and detector response, with associated uncertainties. This model is built by unifying aspects of the two experiments’ analyses where appropriate, and using each individual experiment’s approach otherwise.
Due to the similarities in neutrino energy spectra and event selections, T2K and low energy atmospheric events are described by a common neutrino interaction model.
The remaining SK atmospheric events correspond to neutrinos of higher energies, for which the constraint on interaction uncertainties coming from the lower energy neutrino measurements at the T2K ND is not always directly applicable.
As a result, for these events, the same base interaction model is used for the initial simulation, but largely independent parameters are used for interaction systematic uncertainties.
The neutrino flux models of the two experiments [8, 10, 11] are mostly independent, with the only common source of systematic uncertainty coming from hadron production in proton collisions. Hadron production is tuned using different measurements in the two models: the
SK atmospheric flux model uses atmospheric muon measurements [12, 13], whereas T2K’s flux model uses measurements by the NA61/SHINE experiment [14]. The resulting uncertainties on the flux prediction are therefore considered to be independent between the two models.
The neutrino events of the two experiments are observed in the same detector, and the correlated effects of detector systematic uncertainties on SK and T2K event samples were evaluated for the joint analysis.
Event selection—This analysis uses a total of eighteen SK atmospheric and five T2K event samples, constructed as described in Refs. [9] and [15]. The event selections are based primarily on the number of reconstructed Cherenkov rings, the type of those rings, and the number of delayed Michel electron candidates. The ring types are either showering (-like) or non-showering (-like) and are the basis of the separation between and events.
The T2K selections target events with little activity in the SK outer detector, so-called fully contained (FC) events, with a single Cherenkov ring.
This topology primarily selects charged current quasi-elastic (CCQE)-like events, although in neutrino running mode an additional sample probes events containing a below-Cherenkov-threshold by requiring exactly one -like ring and one Michel electron.
Atmospheric neutrinos span a much wider range of energies, and
the atmospheric FC sample is divided into sub- and multi-GeV categories based on the deposited visible energy in the detector.
The SK analysis additionally includes events with significant energy deposition in the outer detector.
The T2K beam samples and FC single-ring sub-GeV atmospheric samples in SK have a large kinematic overlap, but differ slightly in their respective event selections, which remain the same as in the publications above.
One additional selection criterion, however, is applied to both the SK FC and T2K samples to remove neutron contamination from the Michel candidates for each event, which changes the event rates by .
Interaction model—Neutrino interactions are simulated with the NEUT generator v5.4.0 [16] using the same configuration as T2K’s analysis [8]. The common “low-energy” uncertainty model used for the T2K and atmospheric SK FC sub-GeV samples is based on T2K’s model, with additions to cover important uncertainties for the SK atmospheric analysis. Additional normalization uncertainties on the neutral current single model are introduced, which scale the resonant and coherent contributions separately, motivated by studies of MiniBooNE data [17, 18].
A supplementary uncertainty on the CCQE cross-section ratio is added based on the difference of this ratio between the spectral function model [19] used in this analysis and new calculations using the Hartree-Fock model with Continuum Random Phase Approximation [20, 21].
The ability of the low energy model to describe the atmospheric sub-GeV data samples is evaluated by comparing its predictions using T2K’s ND constraint to the observed data for the down-going part of those samples. The down-going events are mostly unaffected by oscillations, and can therefore be used to test the model without biasing the oscillation measurement.
Good agreement with data is found for the samples targeting events without pions in the final state. For the samples targeting charged current single charged pion (CC1) events, a significant data excess is seen: 225 events are observed for 160.0 12.6(stat) 14.3(syst) predicted in the -like sample, and 84 are observed in the -like sample for 52.0 7.2(stat) 5.3(syst) predicted. The excess is localized at low lepton momentum in the -like sample, and uniformly distributed in the -like sample. An excess is also seen at low momentum in the corresponding -like beam sample, but is not significant due to the low statistics of this sample.
As a result of the observed excesses, an interaction model uncertainty is added to change the shape of the pion three-momentum spectrum for charged-current resonant interactions by modifying the Adler angle [22] distribution, based on theoretically motivated [23, 24] and empirical modifications. A reconstruction uncertainty is also added, as detailed in the next section.
The model for the remaining SK samples (“high-energy model”) is based on that used in the SK analysis, but shares the CCQE part with the low-energy model with the exception of the high (four-momentum transfer) parameters, due to similarities in the CCQE events’ phase space outside of the high region.
SK and T2K’s uncertainty models for resonant production and deep inelastic scattering events are very similar, the main difference being the existence of additional parameters and different nominal values in SK’s analysis. For consistency, the single pion production parameters are set to the same values as in T2K, which come from studies of external data. This increases uncertainties compared to SK’s analysis. The model for pion final-state interactions is tuned to external data [25], as done by T2K.
Data from T2K’s ND are used to constrain uncertainties on the interaction parameters from the low-energy model and the correlated part of the high-energy model, but not on the new uncertainties added for this analysis, or any other aspect of the high-energy model, due to the lack of overlapping phase space between the near-detector and non-sub-GeV atmospheric samples.
Detector model—The same version of the SK event reconstruction algorithm [9] is used for the two experiments, to be able to derive correlations between their detector and reconstruction systematic uncertainties. Many of the detector uncertainties in both SK and T2K’s analyses are estimated from comparisons between atmospheric data and simulation. For these, correlated uncertainties are constructed by evaluating the effects of variations of the detector parameters of the SK model on the number of events in the different samples of the two experiments simultaneously.
Correlations between the reconstructed momentum scale uncertainty of the two experiments are found to have an impact on the constraint on obtained in the data fit. Three of the four analyses described in the next section treat the corresponding parameters as correlated, while the fourth considers them uncorrelated.
The remaining detector uncertainties from the reference SK and T2K analyses that are relevant for only one of the experiments are applied to the corresponding samples.
An additional systematic uncertainty is introduced for the sub-GeV samples targeting CC events.
It allows single ring single Michel electron events with low lepton momentum to migrate between -like and -like samples. The size of this uncertainty covers the excess in data observed for the down-going CC -like events at low momentum.
Oscillation analysis—Two Bayesian and two frequentist analyses based on the model described above but with differences in technical implementation, binning, and statistical methodology are used to fit the data.
To constrain uncertainties using the T2K ND, one analysis performs simultaneous fits of the T2K ND, T2K far detector and SK atmospheric data, while the others use a covariance matrix to encode T2K’s ND constraint on systematic parameters when fitting the events observed at SK.
The Bayesian analyses use Markov Chain Monte Carlo methods to evaluate marginal likelihoods for the parameters of interest, while the frequentist analyses compute the profile likelihood on a fixed grid of values of the oscillation parameters of interest.
For atmospheric oscillation probability calculations, path-dependent density averaging of the matter effect based on a four layer approximation of the PREM model [26] is used, and the fast atmospheric oscillations at low energy are smeared using different techniques depending on the analysis.
The path-dependence yields more precise oscillation probabilities than the conventional approximation assuming layers of constant density.
All analyses utilize a binned log-likelihood test statistic assuming Poisson-distributed bin contents plus Gaussian penalties for the systematic parameters. Finally, reactor experiment measurements of using disappearance, [2, 27, 28, 29, 30], are used as an external constraint.
Robustness studies—Simulated datasets [8] are used to test the robustness of the analysis to effects not included in the base systematic uncertainty model. They are generated using alternative models, and fitted using the nominal model to measure how the obtained -values and constraints on oscillation parameters would be affected by having assumed the wrong model.
Fourteen such simulated datasets are considered, corresponding to alternative neutrino interaction models and data-driven effects at both T2K ND and SK.
These studies are used to estimate, amongst other things, how much the data excess observed in the atmospheric down-going CC samples could bias the results despite the newly added uncertainties if it was coming from an unknown systematic effect. Some of the simulated datasets produce a visible shift in the preferred values for . The uncertainty on is inflated by to account for these effects.
Dataset— The atmospheric dataset is slightly increased compared to [9] to include the full Super-Kamiokande IV period (2008–2018), corresponding to a total live-time of 3244.4 days. The same T2K dataset as in [8] is used, corresponding to exposures of and protons on target in neutrino and antineutrino running modes, respectively.
Bayesian results—The Bayesian analyses assume uniform priors on or , , , and the MO. They find a preference for the normal ordering, and a weak preference for the upper octant (Table 1).
SK and T2K data prefer different octants which leads the joint analysis to have a weaker octant constraint than the individual experiments and a higher probability for maximal mixing, while for the CP-violating phase, both experiments favor similar values (Figure 1).
The exclusion of the CP conserving values of (Figure 2) and is quantified by determining the largest fraction of the posterior density for which the value considered is not included in either of the corresponding highest posterior density credible intervals obtained by the two Bayesian analyses (Table 2).
Table 1:
Octant and MO posterior probabilities using either the full dataset or samples from only one experiment and assuming equal prior probabilities. Values obtained by the second analysis are shown in parentheses.
SK only
T2K only
SK+T2K
Upper octant
0.318 (0.337)
0.785 (0.761)
0.611 (0.639)
Normal ordering
0.654 (0.633)
0.832 (0.822)
0.900 (0.887)
Figure 1: The (, ) credible regions obtained with the SK, T2K, and combined datasets. The MO is marginalized over and a prior uniform in is used.
Table 2:
Largest credible interval from the Bayesian analyses not containing different CP conserving values of and . Values in parentheses indicate how these could change due to possible biases seen in robustness studies.
Value tested
Prior uniform in
Figure 2: The posterior density for the Jarlskog invariant with credible intervals overlaid, marginalized over both MOs, and assuming a uniform prior in either or .
Frequentist results—Ensembles of pseudo-experiments are constructed
to evaluate the frequentist significance of the CP and MO results, taking into account statistical fluctuations and randomizing the values of nuisance oscillation and systematic parameters according to their posterior [31] and prior probability distributions respectively.
Estimating the significance of CP conservation (CPC) based on the presence or not of both and in the confidence intervals was found to have significant over-coverage.
Instead, the log-likelihood ratio between assuming CPC (, here equivalent to ) and without any assumption is used as a test statistic.
For the neutrino MO, the log-likelihood ratio between normal and inverted ordering is used (Figure 3).
The obtained -values are summarized in Table 3. CPC is disfavored with a lower -value () than when using only the T2K data ().
Good agreement () is found with an ensemble that allows for CP-violation by assuming posterior-distributed values.
The inverted ordering is disfavored while good agreement with the normal ordering hypothesis is found, resulting in a parameter [32] for the inverted ordering of 0.18. The distribution of the MO test statistic depends on the assumed values of (from SK samples) and (from T2K samples). The -value for the inverted ordering varies between 0.05 and 0.08 when assuming different fixed true values for and over the range of their 90% confidence intervals. The best-fit values and 68.3% confidence intervals obtained using the Feldman–Cousins method [33] where necessary, are ,
where MO was treated as a nuisance parameter,
and () () for normal (inverted) ordering.
Table 3:
Frequentist -values for the different CP and MO hypotheses.
The most conservative of the two values obtained by the frequentist analyses is given. “-studies” corresponds to the value up to which each -value could increase due to biases seen in robustness studies.
Hypothesis
-value
-studies
CP conservation
0.037
0.050
Inverted ordering
0.079
0.080
Normal ordering
0.58
Figure 3: Distribution of the MO test statistic under true normal and inverted ordering hypotheses. The filled areas to the left (right) of the data result indicate the -values for the inverted (normal) hypotheses.
Goodness of fit—The Bayesian analyses find good posterior predictive -values [34] using both the event spectra ()
and total event counts ().
The -values for the individual T2K samples agree
with the reference T2K analysis [8] up to small differences coming predominantly from model changes.
The frequentist -values [35] additionally show good consistency between the values of the systematic parameters favored by the T2K ND and atmospheric data (), as well as between the atmospheric and beam samples ().
Discussion—The SK and T2K datasets favor similar values for the CP phase, close to maximal CP violation, and both show a preference for the normal MO. As a result, the combined analysis finds increased preferences for non-conservation of CP symmetry and the normal ordering.
When looking directly at the exclusion of CPC through the presence of in credible intervals or frequentist -values, an exclusion at the 2 level is found. However, the significance can fall below 2 due to the potential weaknesses of the uncertainty model tested using simulated datasets. The alternative models having the most impact are the one assuming that the excess seen for down-going CC1-like events is fully due to an unknown systematic effect, and an alternative nuclear model for CCQE interactions [36]. This indicates areas where the robustness of the model should be improved for future analyses.
Conclusion—The SK and T2K collaborations have produced a first joint analysis of their data. A common neutrino interaction and detector model has been developed for the events of the two experiments which overlap in energy, and is found to be able to properly describe the two datasets. The results show an exclusion of the CP-conserving value of the Jarlskog invariant with a significance between 1.9 and 2.0, a limited preference for the normal ordering, and no strong preference for the octant. This first joint analysis is also an important step towards the combined beam and atmospheric data analyses planned by next generation neutrino oscillation experiments.
Acknowledgements
The Super-Kamiokande collaboration gratefully acknowledges cooperation of the Kamioka Mining and Smelting Company.
The Super-Kamiokande experiment was built and has been operated with funding from the
Japanese Ministry of Education, Science, Sports and Culture,
and the U.S. Department of Energy.
The T2K collaboration would like to thank the J-PARC staff for superb accelerator performance. We thank the CERN NA61/SHINE Collaboration for providing valuable particle production data. We acknowledge the support of MEXT, JSPS KAKENHI and bilateral programs, Japan; NSERC, the NRC, and CFI, Canada; the CEA and CNRS/IN2P3, France; the DFG, Germany; the NKFIH, Hungary; the INFN, Italy; the Ministry of Science and Higher Education (2023/WK/04) and the National Science Centre (UMO-2018/30/E/ST2/00441, UMO-2022/46/E/ST2/00336 and UMO-2021/43/D/ST2/01504), Poland; the RSF (RSF 24-12-00271) and the Ministry of Science and Higher Education, Russia; MICINN and ERDF funds and CERCA program, Spain; the SNSF and SERI, Switzerland; the STFC and UKRI, UK; the DOE, USA; and NAFOSTED (103.99-2023.144), Vietnam. We also thank CERN for the UA1/NOMAD magnet, DESY for the HERA-B magnet mover system, the BC DRI Group, Prairie DRI Group, ACENET, SciNet, and CalculQuebec consortia in the Digital Research Alliance of Canada, GridPP and the Emerald High Performance Computing facility in the United Kingdom, and the CNRS/IN2P3 Computing Center in France. In addition, the participation of individual researchers and institutions has been further supported by funds from the ERC (FP7), “la Caixa” Foundation, the European Union’s Horizon 2020 Research and Innovation Programme under the Marie Sklodowska-Curie
grant; the JSPS, Japan; the Royal Society, UK; French ANR and Sorbonne Université Emergences programmes; the VAST-JSPS (No. QTJP01.02/20-22); and the DOE Early Career programme, USA. For the purposes of open access, the authors have applied a Creative Commons Attribution license to any Author Accepted Manuscript version arising.